# Parameters Inference for Nonlinear Wave Equations with Markovian Switching

### v1

@@ Abstract
Traditional partial differential equations with constant coefficients often struggle to capture abrupt changes in real-world phenomena, leading to the development of variable coefficient PDEs and Markovian switching models. Recently, research has introduced the concept of PDEs with Markov switching models, established their well-posedness and presented numerical methods. However, there has been limited discussion on parameter estimation for the jump coefficients in these models. This paper addresses this gap by focusing on parameter inference for the wave equation with Markovian switching. We propose a Bayesian statistical framework using discrete sparse Bayesian learning to establish its convergence and a uniform error bound. Our method requires fewer assumptions and enables independent parameter inference for each segment by allowing different underlying structures for the parameter estimation problem within each segmented time interval. The effectiveness of our approach is demonstrated through three numerical cases, which involve noisy spatiotemporal data from different wave equations with Markovian switching. The results show strong performance in parameter estimation for variable coefficient PDEs.
@@ 1 Introduction
Partial differential equations(PDEs) are crucial in various fields like image processing[1, 2], medical imaging[3], and machine learning[4, 5, 6] for their ability to model real-world phenomena. In automatic control systems, they describe state changes over time and space, aiding in stability and control design[7, 8]. PDEs also model fault monitoring in distributed systems[9] and simulate chemical diffusion in biological cells[10]. However, most of these systems modeled by fixed coefficient equations often exhibit jumps due to perturbations from environmental changes or interventions, and such perturbations can abruptly alter system structures or parameters[11].
@@
To better describe system changes, variable coefficient PDEs have been proposed with theoretical and practical advancements. For instance, Muhammad Sohail and S. Abbas examined the flow of Prandtl fluid influenced by variable coefficient PDEs, focusing on heat and mass transfer using advanced diffusion and conductivity models[12]. Metcalfe and Tataru constructed outgoing parametrices for time-dependent wave equations[13]. Ning et al. examined the stability of wave equations with variable coefficients and nonlinear boundary feedback[14]. Compared to variable coefficient PDEs, another model called the Markovian switching model introduced by Krasovskii in 1961[15] is more adept at handling abrupt changes in system behavior. Early studies about Markovian switching focused on controllability and stability. Practical applications emerged with Sworder’s work on hybrid linear systems affected by Markov jump parameters, providing optimal control feedback[16]. Markovian switching systems have since been applied in various fields. A notable instance is in economics, where Hamilton utilized a Markovian switching model to represent time series behaviors across different regimes. By allowing transitions between multiple structures, this model effectively captures more complex dynamic patterns[17, 18].
@@
Even though both variable coefficient PDEs and the Markovian switching model can address the jumps in a system’s structure or parameters, they still face challenges in capturing the dynamic changes in space, especially across multiple spatial dimensions. Therefore, Li and Wang et al.[19] recently proposed a novel model called diffusion equation with Markovian switching. This model integrates the spatial descriptive power of PDEs with the stochastic nature of Markovian switching mechanisms, allowing for the simultaneous consideration of spatial distribution and temporal jumps in the system. They provided proof of well-posedness, stability, and numerical generation methods using non-uniform grids, along with error analysis. Diffusion equations effectively model processes involving gradual changes such as heat conduction, while wave equations are suited for dynamic phenomena like sound and electromagnetic waves. Incorporating Markovian switching allows wave equations to account for abrupt changes, expanding their applicability. Thus They also constructed the well-posedness for the wave equation with Markovian switching[20]. On the other hand, the inverse problems of some models mentioned above, i.e. inferring parameters from observed data, have garnered attention. Yuan Ye et.al. proposed a framework discovering underlying PDEs and their parameters by sparse Bayesian learning(SBL)[21]. D.A. Barajas-Solano presents two approximate Bayesian inference methods for estimating parameters in PDE models with space-dependent and state-dependent parameters, focusing on diffusion coefficients in linear and nonlinear diffusion equations[22]. Chen and Miao proposed a variable coefficient physics-informed neural network dealing with the forward and inverse problem of variable coefficients PDEs and analyzed the relationship between the convexity of variable coefficients and learning[23]. Li et.al. estimated parameters of a one-dimensional reaction-diffusion model named Fisher’s equation with Markovian switching where the Markov process represents the diffusion parameter[19].
@@
The paper is structured as follows: Section 2 introduces the definition of the wave equation with Markovian switching. Subsequently, the DSBL algorithm is developed for this system, with a detailed discussion of its convergence and error bound. In Section 3, we generate spatiotemporal data through numerical methods and conduct three numerical experiments of increasing complexity, which demonstrate the algorithm’s accuracy in parameter estimation. Our method significantly outperforms a single model without Markovian switching. Finally, Section 4 offers conclusions and prospects for the future work.
@@ 2.1 Definition of Wave equation with Markovian switching
In the real world, the parameters of complex systems usually vary over time. Such coefficient changes can follow stochastic processes including the Markov process, denoted by $M_{t}$. In this paper, we characterize wave equation with Markovian switching in the following form.
$\begin{cases}u_{tt}-M_{t}\Delta u=f(u),&\textit{in}\ U\times(0,T],\\ u(x,t)=0,&\textit{on}\ \partial U\times[0,T],\\ u(x,0)=g(x),&\textit{on}\ U\times\{t=0\}.\end{cases}$ (2.1)
Here, $g(x)$ is the initial value, and $f(u)$ is a nonlinear function. The global time $T$ is fixed and $U\subseteq\mathbb{R}$ is open and bounded. $M_{t}=\{M(t)|0\leq t\leq T\}$ is a Markov process with a state space $\mathcal{E}$. For any $0\leq t_{1}<t_{2}<...<t_{n}\leq T$, there are corresponding states $m_{1},m_{2},...,m_{n}\in\mathcal{E}$. Therefore, $M_{t}$ represents the parameters of the complex system in each state. In particular, the Markovian property regulates that the current value of the state variable depends on its immediate past value. When the system transitions to a new state, the initial value in the current state is the value taken at the end of the previous state, ensuring the continuity of the system.
@@ 2.2 Parameter inference for the wave equation with Markovian switching
Given the structure of PDEs, the parameter inference problem can be converted to a regression problem and then solved using optimization algorithms. Considering the wave systems with Markovian switching, the general form can be rewritten as:
@@
where $t$ and $x=(x_{1},...,x_{n})^{T}\in\mathbb{R}^{n}$, $G$ is an unknown function that depends on the dynamical variable $u$ and its derivatives, $M_{t}$ is a Markov chain, and $\Theta$ represents a parameter switching space. Note that the Markov chain $M_{t}$ governs the parameter vector set $\vartheta=\{\theta_{1},\theta_{2},...,\theta_{d}\}$, which randomly varies within space $\Theta$. Our aim is to learn $G$ and $\vartheta$, thus given a instantaneously varying parameters vector $\theta\in\vartheta$, Eq.2.2 can be written as the following regression form:
@@
Alternatively, by substituting spatio-temporal data into Eq.2.3, we obtain:
@@
Therefore, various methods can be applied to Eq.2.3. In our paper, we primarily develop the SBL method into a discrete optimization algorithm called DSBL. For each discrete state, the SBL optimization algorithm is used for parameter estimation. The DSBL method retains the ability to automatically determine the model complexity and encourages sparser coefficients, achieving dimensionality reduction by introducing a sparse prior distribution $p(\theta;\gamma)$ ($\gamma$ relates to the variance of each component of $\theta$). Similar to SBL, a difference of convex function with respect to the parameters $\theta$ and $\gamma$ is derived, serving as the loss function $\mathcal{L}(\theta,\gamma)$. Subsequently, an iterative algorithm, the Convex-Concave Procedure (CCP), is utilized to solve the following sequence of convex programs:
@@
where $\Omega:=\mathbb{R}^{m}\times[0,+\infty)$.
@@ 2.3 Convergence and error bound of DSBL using in the wave systems with Markovian switching
After presenting the model and its corresponding algorithm, the next essential step is to address the convergence and error bound of the DSBL algorithm. Prior to that, proving the convergence requires establishing the boundedness of the loss function’s and parameters’ iterative sequence, which ensures the existence of a convergent subsequence for $(\theta^{(k+1)},\gamma^{(k+1)})$.
@@
Assume that in Eq.2.1, $M_{t}$ is either a continuous, irreducible, aperiodic, and recurrent Markov chain with a finite state space. The convergence of DSBL can be converted to SBL on the local time interval $[\tau_{k},\tau_{k+1})$. Then, it holds over $[0,T]$,
@@ 3.1 Case 1: Sine-Gordon Equation with Markovian switching
and admits several types of special solutions, such as traveling-wave solutions, soliton solutions and small amplitude solutions. Our goal is to extract a dynamical regime from the soliton solutions.
@@
Where $\alpha$ and $\omega$ represent the parameters of the system. The initial condition $u(x,0)$ is characterized by a Gaussian-shaped curve. It should be noted that $M_{t}$ represents a continuous Markov chain generating by a generator matrix $Q_{1}$ in Eq.3.5 and taking in the following form as Eq.3.4:
@@ 3.2 Case 2: Klein-Gordon Equation with Markovian switching
Similarly, with the first case, we generate the data by the finite difference method and consider the following initial boundary value problem:
@@
We address the problem Eq. 3.7 following a similar procedure to that used for the SG equation, with the exception of employing different grid partitioning methods. Given the changes in the Markov chain, the simulation time is divided into five segments. Then for each segment, we use a grid size of $512\times 512$ for space and time.
@@ 3.3 Case 3: A two-dimensional nonlinear wave equation with Markovian switching
The preceding two cases are both one-dimensional. Now we consider a two-dimensional nonhomogeneous wave equation with Markovian switching. Moreover, we compare the performance of the mixture model and the single model which refers to the process of applying generated data to traditional parameter estimation methods, such as regularization techniques or the least squares method. Here, the comparison is made with the standard SBL algorithm. The specific form of the equation is given as follows:
@@
We define a function $\mathcal{J}$ that maps the time interval to a finite set of system parameter changes:
@@
$\mathcal{J}$ represents the jump points of the parameters controlled by the Markov chain $M_{t}$ over a given time period. In this case, the system’s jump behavior is given by:
@@
Time Interval 2D Wave Equation(Absolute error
@@
Obviously, the shape of the wave changes over time. After applying the DSBL and SBL algorithm to the spatiotemporal dataset respectively, the parameter inference results are presented in Table. 3. It is evident that the mixture model using DSBL demonstrates high accuracy and good adaptability to the switching of system coefficients.
Additionally, several of solutions over time are visualized in Fig. 6. The first column represents the evolution of the numerical solution $u(x,t)$ of the mixture model in the simulation experiment over time. The second and fourth columns show the inferred solutions of the model parameters using the DSBL algorithm and the standard SBL algorithm, respectively, as they evolve over time. The third and fifth columns display the absolute errors between the inferred solutions and the numerical solutions obtained using these two methods. It can be observed that, at the beginning of the experiment, both methods handle the wave equation data well. However, the absolute error of the SBL method becomes increasingly larger compared to our method, leading to a decline in estimation accuracy. This may be due to the accumulation of estimation biases over a longer period.
@@ 5 Appendix
$Proof:$ Before presenting the proof of the theorem, several assumptions are provided. The columns of $D$ are nonzero (i.e., $D_{i}\neq 0$ for $i=1,\ldots,m$), and if $\theta_{i}$ and $\gamma_{i}$ converge to zero simultaneously, then only the path in the space $(\theta,\gamma)$ of the form
$\theta_{i}=\mathcal{O}(\gamma_{i}),\quad\text{when}\ \gamma_{i}\rightarrow 0^{+},$ (5.1)
will be considered. This ensures the limit of loss function $\mathcal{L}$ exists along the path $(\theta,\gamma)$. Define the set $\Omega=\mathbb{R}^{m}\times[0,\infty)^{m}$ and note that $\mathcal{L}$ is only defined on $\textit{int}(\Omega)$. Therefore, we need to extend it to $\bar{\Omega}=\textit{int}(\Omega)\cup\partial\Omega$:
First, for any point $(\theta,\gamma)\in\partial\Omega$, if $\theta_{i}$ and $\gamma_{i}$ are both zero or if $\gamma_{i}\neq 0$ for any $i\in\{1,2,\cdots,m\}$, the limit of $\mathcal{L}$ at the point $(\theta,\gamma)$ exists and is unique along the path satisfying the condition above. This limit is referred to as the value of the function $\mathcal{L}$ at the point $(\theta,\gamma)$.
Second, if there exists an $i$ such that $\theta_{i}\neq 0$ and $\gamma_{i}=0$, then the limit of the function $\mathcal{L}$ at the point $(\theta,\gamma)$ along the path satisfying the condition above is positive infinity. In this case, we define $\mathcal{L}(\theta,\gamma)=M$, where $M$ is a sufficiently large number.
Clearly, the boundary set can be divided into two disjoint sets, i.e., $\partial\Omega=\Omega_{1}\cup\Omega_{2}$, and we define
$\Omega_{1}:=\left\{(\theta,\gamma)\mid\text{there exists at leastone }i\text{ such that }\theta_{i}\neq 0\text{ and }\gamma_{i}=0\right\},$
$\Omega_{2}:=\partial\Omega\backslash\Omega_{1}.$
@@
^$\widehat{\mathcal{L}}\left(\theta,\gamma;\theta^{(k)},\gamma^{(k)}\right)$ $:=f(\theta,\gamma)+g\left(\theta^{(k)},\gamma^{(k)}\right)+\nabla_{\gamma}g\left(\theta^{(k)},\gamma^{(k)}\right)^{T}(\gamma-\gamma^{(k)})$ (5.5)
$c^{(k)}$ $:=\nabla_{\gamma}\log\left|\sigma^{2}I_{n}+D\Gamma^{(k)}D^{T}\right|$ (5.6)
$\Gamma^{(k)}$ $:=\operatorname{diag}\left(\gamma_{1}^{(k)},\gamma_{2}^{(k)},\cdots,\gamma_{m}^{(k)}\right).$ (5.7)
@@
$\begin{gathered}\theta^{(k+1)}\in\arg\min_{\theta}\left\{\|y-A\theta\|^{2}+2\sigma^{2}\sum_{i}\sqrt{c_{i}^{(k)}}\left|\theta_{i}\right|\right\},\\ \gamma_{i}^{(k+1)}=\frac{\left|\theta_{i}^{(k+1)}\right|}{\sqrt{c_{i}^{(k)}}},\quad i=1,2,\cdots,m.\end{gathered}$ (5.8)
@@
$c^{(k)}$ $=\nabla_{\gamma}\log\left|\sigma^{2}I_{n}+A\Gamma^{(k)}A^{T}\right|$ (5.9)
$=\operatorname{diag}\left[A^{T}\left(\sigma^{2}I_{n}+A\Gamma^{(k)}A^{T}\right)^{-1}A\right].$
@@
^$\leq\widehat{\mathcal{L}}\left(\theta^{(k)},\gamma^{(k)};\theta^{(k)},\gamma^{(k)}\right)$
$=\mathcal{L}\left(\theta^{(k)},\gamma^{(k)}\right).$
@@
For every càdlàg function $\mathcal{I}(t)$ defined on $[0,T]$ and for every $k>0$, there exists a finite partition $\mathcal{P}=0<t_{1}<\cdots<t_{n}=T$ such that:
@@
Next, we define the jump of the sample path $\mathcal{I}(t)$ of $M_{t}$ at $t$ as:
@@
$S=\bigcup_{n\in\mathbb{N}}S_{\frac{1}{n}}=\{t:\Delta\mathcal{I}\neq 0\}$ (5.15)
is also finite over the bounded interval $[0,T]$.
@@
According to Eq.5.3, we define a mapping from a point to a set, formulated as follows:
$\mathcal{G}:\quad\Omega$ $\longrightarrow\sigma(\Omega)$ (5.16)
@@
Let $S$ be the set of generalized fixed points of $\mathcal{G}$. By Theorem 2.1, we assume $\left\{\left(\theta^{(k)},\gamma^{(k)}\right)\right\}_{k=0}^{\infty}$ is an iterative sequence generated by the point-to-set mapping $\mathcal{G}$, starting from the initial point $\left(\theta^{(0)},\gamma^{(0)}\right)\in\operatorname{int}(\Omega)$. Since the sequence $\left\{\theta^{(k)}\right\}_{k=0}^{\infty}$ is bounded, let $m_{\theta}$ and $M_{\theta}$ be the lower and upper bounds of the sequence $\left\{\theta^{(k)}\right\}_{k=0}^{\infty}$, respectively. Therefore, there exists a compact set $C=\left[m_{\theta},M_{\theta}\right]^{m}\times\left[0,M_{\gamma}\right]^{m}\subseteq\Omega$ such that for any $k=0,1,\ldots$, all points $\left(\theta^{(k)},\gamma^{(k)}\right)$ are within this compact set $C$. Thus, there exists the following convergent subsequence,
@@
$\left(\theta^{\left(k_{s_{j}}+1\right)},\gamma^{\left(k_{s_{j}}+1\right)}\right)\longrightarrow\left(\theta^{**},\gamma^{**}\right)\text{, as }j\rightarrow\infty.$ (5.23)

### v2

# Data-driven Parameter Inference for Nonlinear Wave Equations with Markovian Switching

@@ Abstract
Traditional partial differential equations with constant coefficients often struggle to capture abrupt changes in real-world phenomena, leading to the development of variable coefficient PDEs and Markovian switching models. Recently, research has introduced the concept of PDEs with Markov switching models, established their well-posedness, and presented numerical methods. However, there has been limited discussion on parameter estimation for the jump coefficients in these models. This paper addresses this gap by focusing on parameter inference for the wave equation with Markovian switching. We propose a Bayesian statistical framework using discrete sparse Bayesian learning to establish its convergence and a uniform error bound. Our method requires fewer assumptions and enables independent parameter inference for each segment by allowing different underlying structures for the parameter estimation problem within each segmented time interval. The effectiveness of our approach is demonstrated through three numerical cases, which involve noisy spatiotemporal data from different wave equations with Markovian switching. The results show strong performance in parameter estimation for variable coefficient PDEs.
@@ 1 Introduction
Partial differential equations(PDEs) are crucial in various fields like image processing[1, 2, 3], medical imaging[4], and machine learning[5, 6, 7] for their ability to model real-world phenomena. In the automatic control field, they describe state changes over time and space, aiding in stability and control design[8, 9]. PDEs also model fault monitoring in distributed systems[10] and simulate chemical diffusion in biological cells[11]. However, most of these systems modeled by fixed coefficient equations often exhibit jumps due to perturbations from environmental changes or interventions, and such perturbations can abruptly alter system structures or parameters[12, 13].
@@
To better describe system changes, variable coefficient PDEs have been proposed with theoretical and practical advancements. For instance, Muhammad Sohail and S. Abbas examined the flow of Prandtl fluid influenced by variable coefficient PDEs, focusing on heat and mass transfer using advanced diffusion and conductivity models[14]. Metcalfe and Tataru constructed outgoing parametrices for time-dependent wave equations[15]. Yin et.al. derived a variable coefficient $(2+1)$-dimensional extended cKP equation for Rossby waves, introducing a new transformation to obtain exact solutions[16]. Ning et al. examined the stability of wave equations with variable coefficients and nonlinear boundary feedback[17]. Compared to variable coefficient PDEs, the Markovian switching model introduced by Krasovskii in 1961[18] is more adept at handling abrupt changes in system behavior. Early studies about Markovian switching focused on controllability and stability. Practical applications emerged with Sworder’s work on hybrid linear systems affected by Markov jump parameters, providing optimal control feedback[19]. Markovian switching systems have since been applied in various fields. A notable instance is economics, where Hamilton utilized a Markovian switching model to represent time series behaviors across different regimes. By allowing transitions between multiple structures, this model effectively captures more complex dynamic patterns[20, 21].
@@
Even though both variable coefficient PDEs and the Markovian switching model can address the jumps in a system’s structure or parameters, they still face challenges in capturing the dynamic changes in space. Therefore, Li and Wang et al.[22] recently proposed a model called diffusion equation with Markovian switching. This model integrates the spatial descriptive power of PDEs with the stochastic nature of Markovian switching mechanisms, allowing for the simultaneous consideration of spatial distribution and temporal jumps in the system. They provided proof of well-posedness, stability, and numerical generation methods using non-uniform grids. Diffusion equations effectively model processes involving gradual changes such as heat conduction, while wave equations are suited for dynamic phenomena like sound and electromagnetic waves. Incorporating Markovian switching allows wave equations to account for abrupt changes, expanding their applicability. Thus They also constructed the well-posedness for the wave equation with Markovian switching[23]. On the other hand, the inverse problems of some models mentioned above, i.e. inferring parameters from observed data, have garnered attention. Yuan Ye et.al. proposed a framework discovering underlying PDEs and their parameters by sparse Bayesian learning(SBL)[24]. D.A. Barajas-Solano presents two approximate Bayesian inference methods for estimating parameters in PDE models with space-dependent and state-dependent parameters, focusing on diffusion coefficients in linear and nonlinear diffusion equations[25]. Lu H et.al. introduced fully nonlinear free surface Physics-Informed Neural Networks to address nonlinear free surface wave propagation and inverse problems in wave modeling[26]. Xu et.al. proposed a spatiotemporal parallel PINNs framework to efficiently solve inverse problems in fluid mechanics for high Reynolds number turbulent flows[27]. Li et.al. estimated parameters of a one-dimensional reaction-diffusion model named Fisher’s equation with Markovian switching where the Markov process represents the diffusion parameter[22].
@@
The paper is structured as follows: Section 2 introduces the definition of the wave equation with Markovian switching. Subsequently, the DSBL algorithm is developed for this system, with a detailed discussion of its convergence property and the error bound. In Section 3, we generate spatiotemporal data through numerical methods and conduct three numerical experiments of increasing complexity, which demonstrate the algorithm’s accuracy in parameter estimation. Our method significantly outperforms a single model without Markovian switching. Finally, Section 4 offers conclusions and prospects for future work.
@@ 2.1 Definition of Wave equation with Markovian switching
In the real world, the parameters of complex systems usually vary over time. Such coefficient changes can follow stochastic processes including the Markov process, denoted by $M_{t}$. In this paper, we characterize wave equation with Markovian switching in the following form
$\begin{cases}u_{tt}-M_{t}\Delta u=f(u),&\textit{in}\ U\times(0,T],\\ u(x,t)=0,&\textit{on}\ \partial U\times[0,T],\\ u(x,0)=g(x),&\textit{on}\ U\times\{t=0\},\end{cases}$ (2.1)
[…]
@@ 2.2 Parameter inference for the wave equation with Markovian switching
Given the structure of PDEs, the parameter inference problem can be converted to a regression problem and then solved using optimization algorithms. Considering the wave systems with Markovian switching, the general form can be rewritten as
@@
where $t\in[0,+\infty)$, $x=(x_{1},...,x_{n})^{T}\in\mathbb{R}^{n}$, $G$ is an unknown function that depends on the dynamical variable $u$ and its derivatives, $M_{t}$ is a Markov chain, and $\Theta$ represents a parameter switching space. Note that the Markov chain $M_{t}$ governs the parameter vector set $\vartheta=\{\theta_{1},\theta_{2},...,\theta_{d}\}$, which randomly varies within space $\Theta$. Our aim is to learn $G$ and $\vartheta$, thus given a instantaneously varying parameters vector $\theta\in\vartheta$, the Eq.2.2 can be written as the following regression form
@@
Alternatively, by substituting spatiotemporal data into Eq.2.3, we obtain
@@
Therefore, various methods can be applied to Eq.2.3. In our paper, we primarily develop the SBL method into a discrete optimization algorithm called DSBL. For each discrete state, the SBL optimization algorithm is used for parameter estimation. The DSBL method retains the ability to determine the model complexity automatically and encourages sparser coefficients, achieving dimensionality reduction by introducing a sparse prior distribution $p(\theta;\gamma)$ and $\gamma$ relates to the variance of each component of $\theta$. Similar to SBL, a difference of convex function concerning the parameters $\theta$ and $\gamma$ is derived, serving as the loss function $\mathcal{L}(\theta,\gamma)$. Subsequently, an iterative algorithm, the convex-concave rocedure(CCP) condition, is utilized to solve the following sequence of convex programs
@@
where $\Omega=\mathbb{R}^{m}\times[0,+\infty)$.
@@ 2.3 Convergence and error bound of DSBL using in the wave systems with Markovian switching
After presenting the model and its corresponding algorithm, the next essential step is to address the convergence and error bound of the DSBL algorithm. Before that, proving the convergence requires establishing the boundedness of the loss function’s and parameters’ iterative sequence, which ensures the existence of a convergent subsequence for $(\theta^{(k+1)},\gamma^{(k+1)})$.
@@
Assume that in Eq.2.1, $M_{t}$ is a continuous, irreducible, aperiodic, and recurrent Markov chain with a finite state space. The convergence of DSBL can be converted to SBL on the local time interval $[\tau_{k},\tau_{k+1})$. Then, it holds over $[0,T]$,
@@ 3.1 Case 1: Sine-Gordon Equation with Markovian switching
and admits several types of special solutions, such as traveling-wave solutions, soliton solutions, and small amplitude solutions. Our goal is to extract a dynamical regime from the soliton solutions.
@@
where $\alpha$ and $\omega$ represent the parameters of the system. The initial condition $u(x,0)$ is characterized by a Gaussian-shaped curve. It should be noted that $M_{t}$ represents a continuous Markov chain generating by a generator matrix $Q_{1}$ in Eq.3.5 and taking in the following form as Eq.3.4
@@ 3.2 Case 2: Klein-Gordon Equation with Markovian switching
Similarly, with the first case, we generate the data by the finite difference method and consider the following initial boundary value problem
@@
We address the problem Eq. 3.7 following a similar procedure to that used for the SG equation, except employing different grid partitioning methods. Given the changes in the Markov chain, the simulation time is divided into five segments. Then for each segment, we use a grid size of $512\times 512$ for space and time.
@@ 3.3 Case 3: A two-dimensional nonlinear wave equation with Markovian switching
The preceding two cases are both one-dimensional. Now we consider a two-dimensional nonhomogeneous wave equation with Markovian switching. Moreover, we compare the performance of the mixture model and the single model which refers to the process of applying generated data to traditional parameter estimation methods, such as regularization techniques or the least squares method. Here, the comparison is made with the standard SBL algorithm. The specific form of the equation is given as follows
@@
We define a function $\mathcal{J}$ that maps the time interval to a finite set of system parameter changes
@@
$\mathcal{J}$ represents the jump points of the parameters controlled by the Markov chain $M_{t}$ over a given period. In this case, the jump behavior of PDEs’ parameter governed by the Markov chain is given by
@@
Time Interval 2D Wave Equation(Absolute error)
@@
The shape of the wave changes over time. After applying the DSBL and SBL algorithm to the spatiotemporal dataset respectively, the parameter inference results are presented in Table. 3. The mixture model using DSBL demonstrates high accuracy and good adaptability to the switching of system coefficients.
Additionally, several solutions over time are visualized in Fig. 6. The first column represents the evolution of the numerical solution $u(x,t)$ of the mixture model in the simulation experiment over time. The second and fourth columns show the inferred solutions of the model parameters using the DSBL algorithm and the standard SBL algorithm, respectively, as they evolve over time. The third and fifth columns display the absolute errors between the inferred solutions and the numerical solutions obtained using these two methods. It can be observed that, at the beginning of the experiment, both methods handle the wave equation data well. However, the absolute error of the SBL method becomes increasingly larger compared to our method, leading to a decline in estimation accuracy. This may be due to the accumulation of estimation biases over a longer period.
@@ 4 Conclusion and future work
Acknowledgements
This work was supported by the NSFC grant 12371198 and the fundamental research funds for the central universities under grand YCJJ20242224.
@@ 5 Appendix
$Proof:$ Before presenting the proof of the theorem, several assumptions are provided. The columns of $D$ are nonzero (i.e., $D_{i}\neq 0$ for $i=1,\ldots,m$), and if $\theta_{i}$ and $\gamma_{i}$ converge to zero simultaneously, then only the path in the space $(\theta,\gamma)$ of the form as
$\theta_{i}=\mathcal{O}(\gamma_{i}),\quad\text{ as }\gamma_{i}\rightarrow 0^{+},$ (5.1)
which will be considered. This ensures the limit of loss function $\mathcal{L}$ exists along the path $(\theta,\gamma)$. Define the set $\Omega=\mathbb{R}^{m}\times[0,\infty)^{m}$ and note that $\mathcal{L}$ is only defined on $\text{int}(\Omega)$. Therefore, we need to extend it to $\bar{\Omega}=\text{int}(\Omega)\cup\partial\Omega$.
For any point $(\theta,\gamma)\in\partial\Omega$, if $\theta_{i}$ and $\gamma_{i}$ are both zero or if $\gamma_{i}\neq 0$ for any $i\in\{1,2,\cdots,m\}$, the limit of $\mathcal{L}$ at the point $(\theta,\gamma)$ exists and is unique along the path satisfying the condition above. This limit is referred to as the value of the function $\mathcal{L}$ at the point $(\theta,\gamma)$. If there exists an $i$ such that $\theta_{i}\neq 0$ and $\gamma_{i}=0$, then the limit of the function $\mathcal{L}$ at the point $(\theta,\gamma)$ along the path satisfying the condition above is positive infinity. In this case, we define $\mathcal{L}(\theta,\gamma)=M$, where $M$ is a sufficiently large number.
The boundary set can be divided into two disjoint sets, i.e., $\partial\Omega=\Omega_{1}\cup\Omega_{2}$, and we define
$\Omega_{1}:=\left\{(\theta,\gamma)\mid\text{there exists at least one }i\text{such that }\theta_{i}\neq 0\text{ and }\gamma_{i}=0\right\},\quad\Omega_{2}:=\partial\Omega\backslash\Omega_{1}.$ (5.2)
@@
The conditional loss function is
^$\widehat{\mathcal{L}}\left(\theta,\gamma;\theta^{(k)},\gamma^{(k)}\right)=f(\theta,\gamma)+g\left(\theta^{(k)},\gamma^{(k)}\right)+\nabla_{\gamma}g\left(\theta^{(k)},\gamma^{(k)}\right)^{T}(\gamma-\gamma^{(k)}),$ (5.5)
where
$c^{(k)}=\nabla_{\gamma}\log\left|\sigma^{2}I_{n}+D\Gamma^{(k)}D^{T}\right|,\quad\Gamma^{(k)}=\operatorname{diag}\left(\gamma_{1}^{(k)},\gamma_{2}^{(k)},\cdots,\gamma_{m}^{(k)}\right).$ (5.6)
@@
$\theta^{(k+1)}\in\arg\min_{\theta}\left\{\|y-D\theta\|^{2}+2\sigma^{2}\sum_{i}\sqrt{c_{i}^{(k)}}\left|\theta_{i}\right|\right\},\quad\gamma_{i}^{(k+1)}=\frac{\left|\theta_{i}^{(k+1)}\right|}{\sqrt{c_{i}^{(k)}}},\quad i=1,2,\cdots,m.$ (5.8)
@@
$c^{(k)}=\operatorname{diag}\left[D^{T}\left(\sigma^{2}I_{n}+D\Gamma^{(k)}D^{T}\right)^{-1}D\right].$ (5.9)
@@
$\leq$ ^$\widehat{\mathcal{L}}\left(\theta^{(k)},\gamma^{(k)};\theta^{(k)},\gamma^{(k)}\right)=\mathcal{L}\left(\theta^{(k)},\gamma^{(k)}\right).$
@@
For every càdlàg function $\mathcal{I}(t)$ defined on $[0,T]$ and for every $k>0$, there exists a finite partition $\mathcal{P}=0<t_{1}<\cdots<t_{n}=T$ such that
@@
Next, we define the jump of the sample path $\mathcal{I}(t)$ of $M_{t}$ at $t$ as
@@
$S=\bigcup_{n\in\mathbb{N}}S_{\frac{1}{n}}=\{t:\Delta\mathcal{I}\neq 0\},$ (5.15)
which is also finite over the bounded interval $[0,T]$.
@@
According to Eq.5.4, we define a mapping from a point to a set, formulated as follows
$\mathcal{G}:\quad\Omega$ $\longrightarrow\sigma(\Omega),$ (5.16)
@@
Let $S$ be the set of generalized fixed points of $\mathcal{G}$. By Theorem 2.1, we assume $\left\{\left(\theta^{(k)},\gamma^{(k)}\right)\right\}_{k=0}^{\infty}$ is an iterative sequence generated by the point-to-set mapping $\mathcal{G}$, starting from the initial point $\left(\theta^{(0)},\gamma^{(0)}\right)\in\operatorname{int}(\Omega)$. Since the sequence $\left\{\theta^{(k)}\right\}_{k=0}^{\infty}$ is bounded, let $m_{\theta}$ and $M_{\theta}$ be the lower and upper bounds of the sequence $\left\{\theta^{(k)}\right\}_{k=0}^{\infty}$, respectively. Therefore, there exists a compact set $C=\left[m_{\theta},M_{\theta}\right]^{m}\times\left[0,M_{\gamma}\right]^{m}\subseteq\Omega$ such that for any $k=0,1,\ldots$, all points $\left(\theta^{(k)},\gamma^{(k)}\right)$ are within this compact set $C$. Thus, there exists the following convergent subsequence
@@
$\left(\theta^{\left(k_{s_{j}}+1\right)},\gamma^{\left(k_{s_{j}}+1\right)}\right)\longrightarrow\left(\theta^{**},\gamma^{**}\right)\quad\text{, as }j\rightarrow\infty.$ (5.23)